# Sociodynamics-inspired Adaptive Coalition and Client Selection in Federated Learning

### v1

@@ Abstract
Federated Learning (FL) enables privacy-preserving collaborative model training, yet its practical strength is often undermined by client data heterogeneity, which severely degrades model performance. This paper proposes that data heterogeneity across clients’ distributions can be effectively addressed by adopting an approach inspired by opinion dynamics over temporal social networks. We introduce FedCVR-Bolt (Federated Coalition Variance Reduction with Boltzmann Exploration), a variance-reducing selection algorithm in which (1) clients dynamically organize into non-overlapping clusters based on asymptotic agreements, and (2) from each cluster, one client is selected to minimize the expected variance of its model update. Our experiments show that in heterogeneous scenarios our algorithm outperforms existing FL algorithms, yielding more accurate results and faster convergence, validating the efficacy of our approach.
@@ 1 Introduction
Federated Learning (FL) enables collaborative model training across multiple clients without sharing raw, potentially sensitive data mcmahan2017communication . While FL offers advantages in privacy, communication efficiency, and access to diverse data, its effectiveness is often limited by statistical heterogeneity. In practice, client data is typically Non-IID, with each local dataset following a distinct distribution zhao2018federated ; li2020federated . This distribution shift leads to inconsistent local updates, slower convergence, and reduced generalization, ultimately compromising the benefits of collaboration.
@@
To address statistical heterogeneity, several approaches have been proposed. Regularization methods constrain local updates to remain close to the global model li2020federated ; li2019fedmd ; karimireddy2020scaffold ; karimireddy2020mime , while personalization techniques adapt models to individual clients or client groups smith2017federated ; finn2017model . Clustering-based methods further mitigate heterogeneity by grouping clients with similar data distributions, enabling the training of specialized models or more informed aggregation ghosh2020efficient ; sattler2020clustered .
@@
Concurrently, client sampling strategies have gained prominence. In practical deployments, factors such as intermittent client availability, energy constraints, and limited bandwidth necessitate selecting only a subset of clients in each round bonawitz2019towards . While uniform random sampling is a common baseline, growing evidence suggests that intelligent, biased selection can accelerate convergence and improve model quality, especially under heterogeneity cho2022towards ; wang2020federated ; goetz2019active . Complementing these empirical advances, theoretical works donahue2021model ; donahue2021optimality have framed FL as a model-sharing game, analyzing incentive compatibility and the Price of Anarchy under various client behaviors and aggregation strategies. This line of research has inspired our analysis, motivating a reformulation of the client interaction process through the lens of opinion dynamics.
In this paper, we introduce a new in addressing statistical heterogeneity by drawing a novel parallel between the iterative model refinement process in FL and the evolution of beliefs in opinion dynamics FriedProsk2019 ; Mas2010_opinion_exp ; raineri2025_ECC_ext . We conceptualize each client’s model parameters as an opinion that evolves via local training and influences the shared global model. Thus, we propose FedCVR-Bolt (Federated Coalition Variance Reduction with Boltzmann exploration) a client sampling algorithm that actively mitigates heterogeneity. Our method proceeds in two stages: (1) an adaptive coalition detection phase that dynamically groups clients with similar model states—reflecting shared data characteristics—and (2) and sampling one client for each cluster using a biased Boltzmann-like probability measure gallavotti2013statistical that allows to maximize the selection probability of the client that maximizes the intra-coalition variance reduction.
@@
After providing a general overview of the main related work in the field, we introduce in Section 2.1 a formal definition of the standard FL setting considered in this study. We then shift our focus in Section 2.2 to the parallelism with Opinion Dynamics, which serves as the foundation for the proposed method. A detailed presentation of the FedCVR-Bolt Algorithm is given in Section 3, including a step-by-step description and the corresponding pseudocode. The core theoretical results that support the methodology are presented in Section 3.1, while full proofs and additional theoretical considerations are deferred to the Appendix. Finally, in Section 4, we validate FedCVR-Bolt across a range of heterogeneous settings and common benchmarks in FL caldas2018leaf , demonstrating its capability to outperform existing baselines.
Contributions:
  • We propose FedCVR-Bolt , a client sampling algorithm grounded in opinion dynamics, which improves performance in heterogeneous FL settings.
  • We formalize a novel analogy between client model updates and opinion formation, guiding our algorithm’s design.
  • We introduce a synthetic dataset for benchmarking federated regression under controlled heterogeneity.
  • We empirically validate our approach on synthetic and real-world data, demonstrating strong performance under Non-IID conditions
@@
Managing statistical heterogeneity across client data distributions can significantly impede the performance and convergence of the global model. Regularization methods, such as FedProx li2020federated , FedMD li2019fedmd , Scaffold karimireddy2020scaffold , and Mime karimireddy2020mime , aim to constrain local client updates or modify the global aggregation process to improve robustness and convergence in heterogeneous settings. Concurrently, personalization techniques focus on adapting the global model or learning client-specific models tailored to individual data distributions, with notable examples including Per-FedAvg fallah2020personalized and pFedMe t2020personalized . Collaboration and clustering strategies seek to identify and leverage similarities among clients by grouping them, thereby enabling more effective federated training within these identified clusters; prominent approaches in this domain include IFCA ghosh2020efficient , CFL sattler2020clustered , FeSEMlong2023multi . More recent study on clustered FL, show the possibility of building groups through consensus based optimization carrillo2024fedcbo , or via training pairwise discriminator to estimate similarities bao2023optimizing .
Client selection strategies have proven to be an effective approach in FL for handling client heterogeneity. Specific heuristic methods have addressed data heterogeneity by prioritizing clients exhibiting higher local loss on the current global model estimate cho2022towards or by aiming for diverse client participation balakrishnan2022diverse . Other methods have centered on clients’ varying training processes diao2020heterofl , data or update quality liao2024quality , or energy efficiency considerations li2019smartpc . Other methods, as FedCBSzhang2023fedcbs ,prioritize clients in order to reduce sampled class imbalance. More advanced policies, such as Oort lai2021oort and Harmony tian2022harmony , utilize analytical strategies to address the diverse nature of device selection. Other approaches have leaned towards learning-based selection policies, based on reinforcement mechanisms. In these methods, the client selection problem is typically formulated as a Markov Decision Process powell2021reinforcement . Notable examples include FedRank tian2024fedrank , AutoFL kim2021autofl , Favor wang2020optimizing , and FedMarl zhang2022multi .
@@
Game theory provides a rigorous framework for modeling incentives and collaboration in FL. Early work by guazzone2013distributed applied coalition formation games to energy-aware resource management in distributed systems, laying the groundwork for strategic collaboration, though not addressing FL statistical challenges. In FL settings, donahue2021model used Hedonic Games to analyze coalition stability in federated linear regression, deriving MSE-optimal aggregation schemes, but limited their focus to linear models and statistical analysis. This was extended in donahue2021optimality , which framed FedAvg as a coalition game, introducing the Price of Anarchy to quantify inefficiencies in naive collaboration. While they proposed an optimal coalition formation algorithm, its scalability remained limited, though the cost function’s submodularity suggested potential for efficient approximations. blum2021one studied equilibria in different FL collaboration structures (one-for-one vs. all-for-all), characterizing when each reaches optimality. Their analysis, however, relied on stylized utility models. To address dynamic interactions, ota2022coalitional proposed Graphical Coalitional Games, forming coalitions based on synergy measures such as cosine similarity or Improvement Classification Accuracy, with robustness analyses under adversarial settings. While structurally innovative, the game-theoretic modeling was relatively shallow. For incentive-aligned client selection, nagalapatti2021game proposed S-FedAvg, using Shapley values to quantify client utility and exclude unhelpful participants, though computational cost and scale were limiting. Lastly, zhang2022game introduced a framework to incentivize high-quality data contributions, directly tackling input data integrity, but faced challenges in distributed quality assessment and efficiency.
@@ 2.1 Federated Learning Problem
The standard FL mcmahan2017communication involves $K$ clients (agents), each holding its own training data. Over $T$ communication rounds, the clients jointly estimate a global vector of $D$ parameters $\theta_{gl}\in\mathbb{R}^{D}$, often simply called the global model. Ideally, the estimate to be found minimizes the global loss:
$\theta_{gl}\in\arg\min_{\theta\in\mathbb{R}^{D}}\mathcal{L}(\theta),$ (1)
@@
$\theta_{gl}(t+1)=\sum_{k=1}^{K}\alpha_{k}\theta_{k}(t+1)\,.$ (2)
This setting is natural when all clients aim to estimate the same $D$-dimensional parameter vector using only their local datasets, typically under the assumption that all clients share the same model architecture.
@@
Partial participation. Since the number of clients $K$ can be very large, the server typically interacts with only a fraction of them in each round. This partial-participation strategy keeps the throughput and energy consumption manageable, mitigates stragglers, and thereby shortens each communication round without compromising convergence mcmahan2017communication ; Luo2024 . Namely, the global model is communicated to the randomly selected subset of clients $\mathcal{P}_{t}\subseteq\{1,\dots,K\}$ whose size $P=|\mathcal{P}_{t}|$ is kept constant across rounds. Only these selected clients compute their local updates $\theta_{k}(t+1)$ and upload them to the server for aggregation. The server updates the global model by aggregating the received estimates:
$\theta_{gl}(t+1)=\sum_{k\in\mathcal{P}_{t}}\tilde{\alpha}_{k}(t)\,\theta_{k}(t+1).$ (3)
@@
Statistical learning setup. We now assume that each client’s training set is a sample drawn from an underlying data-generating distribution. The parameter vector $\theta_{k}\in\mathbb{R}^{D}$ learned by client $k$ (usually, obtained by minimizing its local empirical risk) thus becomes a random variable. Consequently, the vector $\theta_{k}(t)$ produced by client $k$ in communication round $t$ is a realization of that random variable, which the server then incorporates into the global aggregation. We define the random vector $\theta^{d}=(\theta_{1}^{d},\dots,\theta_{K}^{d})^{\top}\in\mathbb{R}^{K}$ whose entries are the $d$-th components of the clients’ parameter vectors, denoting its mean by $\bar{\theta}^{d}=\mathbb{E}\bigl{(}\theta^{d}\bigr{)}$ and its covariance matrix by $C^{d}=\mathbb{C}ov\bigl{(}\theta^{d}\bigr{)}\in\mathbb{R}^{K\times K}.$ Entries of $C^{d}$ quantify client heterogeneity for the $d$-th model parameter: its diagonal elements capture the variance across clients, and the off-diagonal entries measure pairwise similarity between clients’ values. Without loss of generality, we will assume that $C^{d}$ is an invertible matrix. By Equation (2), the components of the global model are scalar random variables $\theta_{gl}^{d}=\alpha^{\top}\theta^{d}$.
@@ 2.2 Problem Setup and Social Dynamics Analogy
Recent studies show that, under data heterogeneity, the choice of the participating-client set $\mathcal{P}_{t}$ is a critical factor for convergence as biased sampling policies can significantly accelerate and stabilize training ghosh2020efficient ; cho2022towards . Motivated by this observation, we introduce a new FL algorithm that adapts insights from social-dynamics models. Specifically, we propose a client-sampling policy that, in heterogeneous federations, trains the global model on a subset of clients that are most representative of the distinct sub-distributions present in the network. Exploiting the analogy with opinion dynamics and social networks of client behavior in distributed learning, our proposed method leverages this parallelism to address two key challenges: (1) how to cluster the client data distributions; (2) how to select the clients to sample within each cluster.
@@
Specifically, as more widely explained in Appendix A, we study our problem through the lens of Opinion Dynamics, where each client is a node in a time-varying network. The set of agents $\mathcal{K}=\{1,\dots,K\}$ refers to the set of clients of the federations and their “opinion” corresponds to the individual client model $\theta_{k}$. The influence matrix $W(t)$ encodes the pairwise influence between clients at time $t$, capturing phenomena such as homophily Mas2010_opinion_exp , i.e., agents are mainly influenced by individuals who hold very similar opinions. This modeling framework is fundamental to identify coalitions across the network, a central challenge studied both in Clustered FL sattler2020clustered and in social community detection Mark2003 ; Mas2010_opinion_exp . In practical scenarios, it is common to find groups of clients with similar data distributions and, consequently, similar local models. Such clients are naturally inclined to collaborate, making coalition identification a key aspect of effective FL system design. This holds in particular since sample all the clients to update the model data is expensive, making thus the exploitation of this extra information essential. Therefore, our idea is to find $P$ groups of co-operating clients, such that their data comes from the same distribution, and then sample the best client for each group according to a proper strategy explained later on.
@@
A common challenge, originally arising in the context of sensor placement problems kempe2008_subsetselection ; kempe2018_ApproxSubmod , and later extended to social applications in raineri2025_ECC_ext ; raineri2023_ifac , is the estimation of the overall network opinion when the true opinions of individuals are unknown and only a subset can be observed. Analogously, here, our goal is to select the optimal subset of clients to sample to best approximate the global model $\theta_{gl}$. Based on the cited literature approach and building on the statistical learning setup previously introduce, we will introduce a novel sampling strategy based on variance reduction techniques.
3 FedCVR-Bolt Algorithm
Here, we will specifically focus on the main algorithm steps which occur at every communication round $t\in[T]$. At the end of this section, we summarize the procedure in Algorithm 1.
Opinion Dynamics Setting

### v2

# Socially Inspired Coalition Formation and Client Selection in Federated Learning

@@ Abstract
Federated Learning (FL) enables privacy-preserving collaborative model training, but its effectiveness is often limited by client data heterogeneity.We introduce a client-selection algorithm that (i) dynamically forms non-overlapping coalitions of clients based on asymptotic agreement and (ii) selects one representative from each coalition to minimize the variance of model updates. Our approach is inspired by social-network modeling, leveraging homophily-based proximity matrices for spectral clustering and techniques for identifying the most the most informative individuals to estimate a group’s aggregate opinion. We provide theoretical convergence guarantees for the algorithm under mild, standard FL assumptions. Finally, we validate our approach by benchmarking it against three strong heterogeneity-aware baselines; the results show higher accuracy and faster convergence, indicating that the framework is both theoretically grounded and effective in practice.
@@ 1 Introduction
Federated Learning (FL) enables collaborative model training across multiple clients without sharing raw, potentially sensitive data (McMahan et al., 2017). While FL reduces the need to centralize data and can improve privacy and regulatory compliance, its effectiveness is often limited by statistical heterogeneity. In practice, client data is typically non-IID, with each local dataset following a distinct distribution (Zhao et al., 2018; Li et al., 2020). This distribution shift leads to inconsistent local updates, slower convergence, and reduced generalization, undermining the benefits of collaboration. To address statistical heterogeneity, several approaches have been proposed. Regularization methods constrain local updates to remain close to the global model (Li et al., 2020; Li & Wang, 2019; Karimireddy et al., 2020b; a), while personalization techniques adapt models to individual clients or client groups (Smith et al., 2017; Finn et al., 2017). Clustering-based methods further mitigate heterogeneity by grouping clients with similar data distributions, enabling the training of specialized models or more informed aggregation (Ghosh et al., 2020; Sattler et al., 2020).
@@
Concurrently, client-sampling strategies have gained prominence. In practical deployments, factors such as intermittent client availability, energy constraints, and limited bandwidth necessitate selecting only a subset of clients in each round (Bonawitz et al., 2019). While uniform random sampling is a common baseline, growing evidence suggests that non-uniform (biased) selection can accelerate convergence and improve model quality, especially under heterogeneity (Cho et al., 2022; Wang et al., 2020b; Goetz et al., 2019). Complementing these empirical advances, theoretical works have framed FL as a model-sharing game (Donahue & Kleinberg, 2021a; b), analyzing incentive compatibility and the Price of Anarchy under various client behaviors and aggregation strategies.
In this work, we introduce FedCVR-Bolt (Federated Coalition Variance Reduction with Boltzmann exploration), a client sampling algorithm that mitigates heterogeneity via two stages: (1) an adaptive coalition detection phase that dynamically groups clients with similar model states and (2) a within-coalition selection step using a biased Boltzmann-like probability measure to maximize the expected intra-coalition variance reduction (Gallavotti, 2013). Our primary focus is conceptual and methodological: we draw an analogy between client selection in FL and homophily-based social selection mechanisms Mäs et al. (2010), and we leverage results form estimation theory (Kay, 1993) to ground the variance–reduction principle at the core of FedCVR-Bolt .
Paper structure
@@
After a brief overview of related work, Section 2 introduces the FL framework and problem setup, drawing parallels with social and sensor networks that inspire our approach. A detailed presentation of the FedCVR-Bolt is given in Section 3, including a step-by-step description and the corresponding pseudocode. The core theoretical results that support the methodology are presented in Section 3.1, while proofs and additional theoretical considerations are deferred to the Appendix. In Section 5, we validate FedCVR-Bolt across a range of heterogeneous settings and common benchmarks in FL (Caldas et al., 2018), showing its capability to outperform existing baselines.
Contribution
Our contributions are as follows: i) We propose FedCVR-Bolt , a client sampling algorithm, which improves performance in heterogeneous FL settings compared to existing sampling strategies. We draw on social-interaction models to introduce homophily-based proximity matrices for spectral clustering and to adapt most-informative-node selection for client choice. ii) We provide theoretical guarantees on the convergence of the FedCVR-Bolt algorithm under mild assumptions. iii) We introduce a synthetic dataset for benchmarking federated regression under controlled heterogeneity. iv) We empirically validate our approach on synthetic and real-world data, demonstrating strong performance under Non-IID conditions.
@@
Statistical heterogeneity across client data distributions can significantly impede the performance and convergence of the global model. Regularization methods, such as FedProx (Li et al., 2020), FedMD (Li & Wang, 2019), Scaffold (Karimireddy et al., 2020b), and Mime (Karimireddy et al., 2020a), constrain local client updates or modify the global aggregation process to improve robustness and convergence in heterogeneous settings. Personalization techniques adapt the global model or learn client-specific models tailored to individual data distributions, with notable examples including Per-FedAvg (Fallah et al., 2020) and pFedMe (T Dinh et al., 2020). Collaboration and clustering strategies identify and leverage similarities among clients by grouping them, enabling more effective federated training within these identified clusters; prominent approaches in this domain include IFCA (Ghosh et al., 2020), CFL (Sattler et al., 2020), FeSEM (Long et al., 2023). More recent studies on clustered FL show the possibility of forming groups via consensus-based optimization (Carrillo et al., 2024) or by training a pairwise discriminator to estimate similarities (Bao et al., 2023).
Client selection strategies prove to be an effective approach in FL for handling client heterogeneity. Various methods address data heterogeneity by prioritizing clients exhibiting higher local loss on the current global model estimate (Cho et al., 2022) or by aiming for diverse client participation (Balakrishnan et al., 2022). Other approaches focus on variability in clients in training processes (Diao et al., 2020), data and/or update quality (Liao et al., 2024), or energy efficiency (Li et al., 2019). Methods such as FedCBS (Zhang et al., 2023) prioritize clients to reduce sampled class imbalance. More advanced policies, such as Oort (Lai et al., 2021) and Harmony (Tian et al., 2022), employ analytical rules address the device-selection heterogeneity. A distinct research thread develops selection policies with reinforcement learning, typically framing client selection as a Markov decision process (Powell, 2021). Notable examples include FedRank (Tian et al., 2024), AutoFL (Kim & Wu, 2021), Favor (Wang et al., 2020a), and FedMarl (Zhang et al., 2022b).
@@
Game theory provides a rigorous framework for modeling incentives and collaboration in FL. Early work in Guazzone et al. (2013) applied coalition formation games to energy-aware resource management in distributed systems, laying the groundwork for strategic collaboration, though not addressing FL’s statistical challenges. In FL settings, Donahue & Kleinberg (2021a) used hedonic games to analyze coalition stability in federated linear regression, deriving MSE-optimal aggregation schemes, but limited their analysis to linear models. This was extended in Donahue & Kleinberg (2021b), which framed FedAvg as a coalition game, introducing the Price of Anarchy to quantify inefficiencies in naive collaboration. They proposed an optimal coalition formation algorithm of limited scalability, although submodularity of the cost function suggests potential for efficient approximations. Blum et al. (2021) studied equilibria in different FL collaboration structures (one-for-one vs. all-for-all), characterizing when each is optimal (under stylized utility models). To address dynamic interactions, Ota et al. (2022) proposed graphical coalitional games, forming coalitions via synergy measures such as cosine similarity or Improvement Classification Accuracy and analyzing robustness under adversarial settings. For incentive-aligned client selection, Nagalapatti & Narayanam (2021) proposed S-FedAvg, using Shapley values to quantify client utility and exclude unhelpful participants, though computational cost limits scalability. Recently, Zhang et al. (2022a) introduced a framework to incentivize high-quality data contributions, directly tackling input-data integrity, but faced challenges in distributed quality assessment and efficiency.
@@ 2.1 Federated Learning Problem
The standard FL (McMahan et al., 2017) involves $K$ clients (agents), each holding its own training data. Over $T$ communication rounds, the clients jointly estimate a global vector of $D$ parameters $\theta_{gl}\in\mathbb{R}^{D}$, often called the global model. Ideally, the estimate to be found minimizes the global loss:
$\theta_{gl}\in\arg\min\nolimits_{\theta\in\mathbb{R}^{D}}\mathcal{L}(\theta),$ (1)
@@
$\theta_{gl}(t+1)=\sum\nolimits_{k=1}^{K}\alpha_{k}\theta_{k}(t+1)\,.$ (2)
This setting is natural when all clients aim to estimate the same $D$-dimensional parameter vector using only their local datasets, under the assumption that all clients share the same model architecture.
@@
Partial participation. Since the number of clients $K$ can be very large, the server typically interacts with only a fraction of them in each round. This partial-participation strategy keeps the throughput and energy consumption manageable, mitigates stragglers, and thereby shortens each communication round without compromising convergence (McMahan et al., 2017; Luo et al., 2024). Namely, the global model is communicated to the randomly selected subset of clients $\mathcal{P}_{t}\subseteq\{1,\dots,K\}$ whose size $P=|\mathcal{P}_{t}|$ is kept constant across rounds. Only these selected clients compute their local updates $\theta_{k}(t+1)$ and upload them to the server for aggregation. The server updates the global model by aggregating the received estimates:
$\theta_{gl}(t+1)=\sum\nolimits_{k\in\mathcal{P}_{t}}\tilde{\alpha}_{k}(t)\,\theta_{k}(t+1).$ (3)
@@
Statistical learning setup. We now assume that each client’s training set is a sample drawn from an underlying data-generating distribution. The parameter vector $\theta_{k}\in\mathbb{R}^{D}$ learned by client $k$ (usually, obtained by minimizing its local empirical risk) thus becomes a random variable. Consequently, the vector $\theta_{k}(t)$ produced by client $k$ in communication round $t$ is a realization of that random variable, which the server then incorporates into the global aggregation. We define the random vector $\theta^{d}=(\theta_{1}^{d},\dots,\theta_{K}^{d})^{\top}\in\mathbb{R}^{K}$ whose entries are the $d$-th components of the clients’ parameter vectors, denoting its mean by $\bar{\theta}^{d}=\mathbb{E}\bigl(\theta^{d}\bigr)$ and its covariance matrix by $C^{d}=\mathbb{C}ov\bigl(\theta^{d}\bigr)\in\mathbb{R}^{K\times K}.$ Entries of $C^{d}$ quantify client heterogeneity for the $d$-th model parameter: its diagonal elements capture the variance across clients, and the off-diagonal entries measure pairwise similarity between clients’ values. Without loss of generality, we will assume that $C^{d}$ is an invertible matrix. By Equation equation 2, the components of the global model are scalar random variables $\theta_{gl}^{d}=\alpha^{\top}\theta^{d}$.
@@ 2.2 Problem Setup and Social Dynamics Analogy
Recent studies show that, under data heterogeneity, the choice of the participating-client set $\mathcal{P}_{t}$ is critical for convergence as biased sampling policies can significantly accelerate and stabilize training (Ghosh et al., 2020; Cho et al., 2022). Building on this, we introduce a FL algorithm inspired by social-dynamics models that sample clients via a two-step procedure. First, we form coalitions whose covariance mirrors client relationships, explicitly encouraging high intra-cluster and low inter-cluster covariance, so that each cluster captures a distinct sub-distribution. Second, within each cluster, we sample clients to train the global model on representatives that best preserve this structure, thereby directly mitigating the effect of heterogeneity. This address two key challenges: i) how to partition the client data distributions; ii) how to select the clients to sample within each cluster.
@@
Specifically, as explained in Appendix B, in our study we consider each client as a node (or agent) in a time-varying network. The set of agents $\mathcal{K}=\{1,\dots,K\}$ refers to the set of clients of the federations and their “opinion” corresponds to the individual client model $\theta_{k}$. The influence matrix $W(t)$ encodes the pairwise influence between clients at time $t$, capturing phenomena such as homophily (Mäs et al., 2010), i.e., agents are mainly influenced by individuals who hold similar opinions. This modeling framework is fundamental to identify coalitions across the network, a key challenge studied both in Clustered FL (Sattler et al., 2020) and in social community detection (Mark, 2003; Mäs et al., 2010). In practical scenarios, it is common to find groups of clients with similar data distributions and, consequently, similar local models. Such clients are naturally inclined to collaborate, so identifying coalitions is central to effective FL design, aligned with well established methods like Clustered FL (Sattler et al., 2020), IFCA (Ghosh et al., 2020) and FeSEM (Long et al., 2023). Accordingly, we partition the federation into $P$ groups of cooperating clients whose data are drawn from the same (or closely related) distribution, and then, at each round, select the best representative from each group using the strategy detailed below.
@@
A common challenge, originally arising in the context of sensor placement problems (Das & Kempe, 2008; 2018) and later extended to social networks (Raineri et al., 2025; 2023) is the estimation of the overall network opinion when only a subset can be observed. Analogously, here, our goal is to select the optimal subset of clients to sample to best approximate the global model $\theta_{gl}$. Based on the cited literature approach and building on the statistical learning setup previously introduce, we will introduce a novel sampling strategy based on variance reduction techniques.
3 FedCVR-Bolt Algorithm Design
Analogy with Opinion Dynamics